OpenStudy (anonymous):

Find a quadratic model for the set of values.(–2, –20), (0, –4), (4, –20)

OpenStudy (amistre64):

set up an equation that fits the information y = a + b(x+2) + cx(x+2) when x=-2, y = -20 -20 = a + b(-2+2) + c(-2)(-2+2) -20 = a when x=0, y = -4 -4 = -20 + b(0+2) + c(0)(0+2) 16 = b(2) ; b=8 when x=4, y = -20 -20 = -20 + 8(4+2) + c(4)(4+2) 0 = 8(6) + c(4)(6) -48 = c24; c = -2 then fill in the general and expand

OpenStudy (amistre64):

y = -20 +8(x+2) -2x(x+2) y = -20 +8x+16 -2x^2 - 4x y = -4 +4x -2x^2

hartnn (hartnn):

can u tell,why did u take,x+2 instead of y=a+bx+cx*x

OpenStudy (amistre64):

becasue x=-2 is needs to zero out my setup

OpenStudy (amistre64):

i could have used any combination of the 3 x values to zero it out with, but i just used it in order that it was presented

hartnn (hartnn):

ok :) i was wondering why that did not strike me...

OpenStudy (anonymous):

this is the choices i have.A. y = 4x2 + 4x + 2 B. y = 2x2 - 4x + 4 C. y = 4x2 - 2x - 4 D. y = -2x2 + 4x - 4

OpenStudy (amistre64):

and what a nice set of choices they are .....

OpenStudy (amistre64):

if you had choices available to you, then you could have simply plugged in your x and y values to see what choice fit

hartnn (hartnn):

^ exactly...much more simpler

OpenStudy (amistre64):

not as fun, but simpler nonetheless lol